# A Theory of Recommendations

### v1

@@ 5.1 Characterizing the Value of the Recommendation System
$V_{0}(i):=U_{i}^{0}.$ (10)
Now suppose one of the products comes with a buy or don’t-buy recommendation, so the receiver has to update his belief about the quality distribution. A receiver who accepts the recommendation chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation. With a recommendation system, the expected payoff of a receiver $i$ who accepts a recommendation is thus
$V_{A}(i):=\pi^{B}U_{i}^{B}+(1-\pi^{B})U_{i}^{0}.$ (11)
The first term is the probability of receiving a buy recommendation multiplied by the expected payoff from accepting it (and thus buying the recommended product). The second term is the probability of a don’t-buy recommendation multiplied by the expected payoff from buying the product without any recommendation, that is, based on priors. [^15] The notation $V_{A}(i)$ emphasizes the dependence of this expected payoff on the type of the receiver. Clearly, $V_{A}$ also depends on $R$ and on the decision environment; similarly for the expressions below.
A receiver who does not accept a recommendation will buy the recommended product if it comes with a don’t-buy recommendation. Otherwise, he will buy one of the other products. Hence, the expected payoff of such a receiver is
$V_{N}(i):=$ $\pi^{B}U_{i}^{0}+(1-\pi^{B})U_{i}^{D}.$ (12)
A buy recommendation does not create value for receivers who do not follow it, as such types choose the alternative for which no recommendation is available. Instead, the don’t-buy recommendation, which is given with probability $(1-\pi^{B})$, does influence the behavior of the types who do not accept it: Instead of buying an alternative based on prior probabilities, updating leads them to choose the product for which they received the explicit recommendation not to buy.
@@
Intuitively, when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, Corollary 1 implies that all receivers accept the recommendation as the objective effect dominates. Hence, by equation (11), the recommendation system creates value only in case of a buy recommendation, as a don’t-buy recommendation leads to the same expected payoff as without a recommendation system. The positive payoff effect of the recommendation reflects the type-independent objective part $\Delta_{O}^{B}$ and the subjective part $\Delta_{S}^{B}$, which is weighted by the average type. Finally, all of that only materializes with the probability of a buy recommendation, $\pi_{B}$.
@@
As the absolute value of $\Delta_{S}^{B}$ increases, the analysis becomes more complex because some types do not follow the recommendation. We consider the case $\Delta_{S}^{B}>2\Delta_{O}^{B}$ and reproduce the expression for $V(\mathcal{R})$ as follows: [^17] The argument for $\Delta_{S}^{B}<-2\Delta_{O}^{B}$ is analogous.
@@
$\underbrace{\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]}_{\text{types accept recommendation}}+\underbrace{(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]}_{\text{types reject recommendation}}.$ (13)
@@
In this case, the buy recommendation shifts probability from $(1,0)$ to $(0,1)$, which is unfavorable for receivers $i>0$, reducing their gains from accepting and possibly making them negative. Thus, for $\Delta_{S}^{B}>2\Delta_{O}^{B}$, only types below $\tilde{i}$ (a fraction $F(\tilde{i})$ of the receivers) accept the buy recommendation, which comes with probability $\pi^{B}$. In addition to the objective value of the recommendation ($\Delta_{O}^{B}$), these receivers obtain a subjective value. In (13), the subjective value is reflected in $-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]$, the average effect of the shift in the relative probabilities of $(1,0)$ and $(0,1)$ on the value of accepting the recommendation for these buyers. The sign of this term depends on the properties of $F$ captured in $\mathbb{E}[i\mid i\leq\tilde{i}]$. It is positive if the negative contribution of types $i<0$ dominates the positive contribution of types $i>0$. The second component in (13), corresponding to the types who do not accept the recommendation, is arguably more surprising. Given their lack of alignment with the rest of the population, these types treat the recommendation not to buy as good news about the product and thus do not accept the recommendation. [^18] In detail, a don’t-buy recommendation (which happens with probability $1-\pi^{B}$) carries an adverse objective content (summarized by $\Delta_{O}^{D}$ in the term in squared brackets). However, for receivers with $i>0$ it also carries the good news that the product is more likely to be of the preferred version $(1,0)$ rather than $(0,1)$. Those with $i>\tilde{i}(>0)$ then buy the product despite the adverse objective content, valuing the beneficial subjective effect with $-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]>0$ on average. Such receivers exist only if $\Delta_{S}^{B}$ is sufficiently large (and hence $\Delta_{S}^{D}$ is sufficiently negative relative to $\Delta_{O}^{D}$). Nonetheless, its informational content creates value for them.
@@
In the following, we shall often invoke the following symmetry assumption on the preference distribution.
@@
With this assumption in place, the quantity
$\beta:=F(1/2-R)$ (14)
has a simple interpretation. Given a threshold $R$, the parameter $\beta$ captures the fraction of the population that is willing to recommend product versions $(1,0)$ or, equivalently (given the symmetry of the distribution), $(0,1)$. For a fixed type distribution $F$, the variables $\beta$ and $R$ are inversely related. We then obtain the following result:
@@
Suppose Assumption 1 holds. Then, all types accept the recommendation and
@@
To understand why all types accept a recommendation, the belief-updating decomposition in Section 4.1 is helpful. Its third component captures the potential change in the odds of the controversial products, which induces a purely type-dependent payoff effect. Given the symmetry assumption, both controversial product versions lead to a buy recommendation with probability $\beta$. Hence, a recommendation does not change the odds of the controversial products. As the other two effects of a buy recommendation on payoffs are positive for all types, the buy recommendation increases expected payoffs for any type, yielding the result. [^19] An analogous argument applies for don’t-buy recommendations. Proposition 3 thus implies $V(R)=\pi^{B}\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]\right]$. Further, because of the symmetry of the population, $\mathbb{E}[i]=0$, so $V(R)=\pi^{B}\Delta_{O}^{B}$ as required.
@@ 5.2.1 Symmetric Type Distribution
By Lemma 2, an increase in $R$ reduces the probability of a buy recommendation, while it increases $\Delta_{O}^{B}$, the expected average payoff conditional on a buy recommendation. These effects exactly offset each other if and only if $\sigma=1$. Otherwise, one effect dominates the other. Based purely on uncontroversial products, a buy recommendation is more likely than a don’t-buy recommendation when the good product is relatively more likely than the bad one ($\sigma>1$). It is thus worthwhile to increase the expected average payoff conditional on a buy recommendation by increasing $R$ at the expense of the probability of receiving such a recommendation. Put differently, the designer chooses between revealing whether the product is good or whether it is bad. He moves the threshold in the direction which yields full revelation with higher probability, which hinges only on the odds ratio of the uncontroversial product versions.
@@ 5.2.2 Asymmetric Type Distribution
We can now analyze the extent to which the result in Proposition 5 that, generically, extreme thresholds are desirable is an artifact of the symmetry assumption. This is challenging because, contrary to the symmetric case, the set of types that accept a recommendation depends on the threshold $R$. Reflecting this complication, there is no direct counterpart to Proposition 5. Nevertheless, the following two results provide some insights into the optimal recommendation system in the case of an asymmetrically distributed population.
@@
As illustrated in Figure 2, Proposition 6 provides sufficient conditions for interior solutions. It shows that an interior optimum exists when (i) controversial products are likely ($Q$ is large) and the population is very asymmetric ($a$ is far away from $1$) or otherwise when (ii) the uncontroversial product versions are similarly likely ($\sigma$ is close to $1$). The result starkly contrasts the symmetric case, where an interior solution is only possible in the knife-edge case $\sigma=1$. In that case, recommendations do not reveal any information about the relative likelihood of the controversial products beyond the prior, so the subjective payoff effect cancels out and the optimal threshold depends only on the odds ratio $\sigma$ of the uncontroversial products. In contrast, the average subjective payoff effect does not cancel out in the asymmetric case. Thus, the choice of the optimal threshold no longer only depends on $\sigma$ but also on the asymmetry parameter $a$ and the prevalence of controversial products $Q$. For a wide range of parameters, the conditions given in Proposition 6 are necessary as well as sufficient:
@@
The proof essentially proceeds in two steps. First, all types accept the recommendation when $Q$ is sufficiently small, since then the subjective content in the recommendation vanishes.242424 $Q\rightarrow 0$ implies that $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, ensuring that all types follow by Corollary 1. The key step is that $\Delta_{S}^{B}\rightarrow 0$ as $Q\rightarrow 0$, while $\Delta_{O}^{B}\rightarrow 1-q_{H}>0$ when $Q\rightarrow 0$. The proof shows that, even for larger values of $Q$, $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$ often holds, explaining why the conclusion of Proposition 7 holds not only in the limit. This simplifies $V(R)$ by eliminating the case distinctions. Leveraging this, we can prove the quasiconcavity of the function, ensuring that the sufficient conditions in Proposition 6 are necessary, too.
@@
The requirement that $Q\rightarrow 0$ is much stronger than necessary for this result. Even for higher values of $Q$, Figure 2 quite accurately separates the parameter range under which an interior solution exists from the area where it does not. [^25] The figure is drawn for $Q=0.1$, in which case the if-and-only-if statement is correct for $\sigma\in[1/60,60]$ and $a\in[1/10000,10000]$. In a nutshell, the result shows that an interior solution exists if (i) controversial products are sufficiently prevalent, (ii) the population asymmetry is substantial, and (iii) the odds ratio of the uncontroversial products is sufficiently small or large. To understand this intuitively and to contrast it with the symmetric case, suppose that $\sigma>1$. Hence, the value of the recommendation system is increasing in $R$ in the symmetric case, as the gain in the expected payoff conditional on receiving a buy recommendation outweighs the loss from the lower probability of a buy recommendation. This follows because the subjective effect in $V(R)=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$ cancels out with symmetry. With an asymmetric population, the term $-\Delta_{S}^{B}\mathbb{E}[i]$ does not vanish. To illustrate this, consider a very asymmetric preference distribution such as $F(x)=(x+1/2)^{10}$, for which the probability of a type $i\leq 0$ is less than 0.1%. Then, $\mathbb{E}[i]>0$ and, by Corollary 2, $\Delta_{S}^{B}<0$ so that $-\Delta_{S}^{B}\mathbb{E}[i]>0$. Hence, the subjective effect is positive on average. Intuitively, though this population is distributed very asymmetrically around zero, it is essentially homogeneous in the sense that almost all types agree on the ranking of the controversial products: If such a product results in a buy recommendation, it almost surely has quality vector (1,0), as almost all types prefer this to $(0,1)$. As such, these products are not that controversial anymore, and even the subjective effect is “fairly objective.” Now, as $R$ increases, the probability of a buy recommendation still falls, and the “purely objective” payoff part $\Delta_{O}^{B}$ increases. However, for $R$ high enough, this “fairly objective” payoff part $-\Delta_{S}^{B}\mathbb{E}[i]$ also decreases, as an extreme threshold $R$ impedes the ability to distinguish between a good product and a widely preferred controversial product. While consumers are homogeneous regarding which controversial product version they prefer, they are still heterogeneous with respect to the intensity with which they do so. Hence, disentangling the good and the preferred controversial product is worthwhile. Therefore, an interior optimal threshold results.
@@ 5.3.1 Preference Heterogeneity
Under Assumption 1, we can analyze increases in the polarization of the distribution of preferences by keeping the threshold $R$ fixed and varying $\beta$ [^26] For $R=1/2$, we must have $\beta=1/2$ because the symmetry of $F$ requires $F(0)=1/2$, so this exercise is only meaningful for thresholds $R\neq 1/2$. Further, for $R>1/2$ we must have $\beta\leq 1/2$ because $\beta=F(1/2-R)\leq F(0)=1/2$. Accordingly, for $R<1/2$ we must have $\beta\geq 1/2$. This corresponds to a mean-preserving spread of $F$ (see Figure 3), coinciding with an increase in $\beta$ for $R>1/2$ and a decrease for $R<1/2$. We obtain the following result.

### v2

@@ 5.1 Characterizing the Value of the Recommendation System
$V^{0}(i):=U^{0}(i).$ (5)
Now suppose one of the products comes with a buy or don’t-buy recommendation, so the receiver has to update his belief about the quality distribution. A receiver who accepts all recommendations chooses this product (rather than one of the alternatives) only if it comes with a buy recommendation. With a recommendation system, the expected payoff of a receiver $i$ who accepts all recommendations is thus
$V^{A}(i):=\pi^{B}U^{B}(i)+(1-\pi^{B})U^{0}(i).$ (6)
The first term is the probability of receiving a buy recommendation multiplied by the expected payoff from accepting it (and thus buying the recommended product). The second term is the probability of a don’t-buy recommendation multiplied by the expected payoff from buying the product without any recommendation, that is, based on priors. [^16] The notation $V^{A}(i)$ emphasizes the dependence of this expected payoff on the type of the receiver. Clearly, $V_{A}$ also depends on $R$ and on the decision environment; similarly for the expressions below.
A receiver who never accepts recommendations will buy the recommended product if it comes with a don’t-buy recommendation. Otherwise, he will buy one of the other products. Hence, the expected payoff of such a receiver is
$V^{N}(i):=$ $\pi^{B}U^{0}(i)+(1-\pi^{B})U^{D}(i).$ (7)
A buy recommendation does not create value for receivers who do not accept it, as such types choose the alternative for which no recommendation is available. Instead, the don’t-buy recommendation, which is given with probability $(1-\pi^{B})$, does influence the behavior of the types who do not accept it: Instead of buying an alternative based on prior probabilities, updating leads them to choose the product for which they received the explicit recommendation not to buy.
@@
Intuitively, when $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, Corollary 1 implies that all receivers accept all recommendations as the objective effect dominates. Hence, by equation (6), the recommendation system creates value only in case of a buy recommendation, as a don’t-buy recommendation leads to the same expected payoff as without a recommendation system. The positive payoff effect of the recommendation reflects the type-independent objective part $\Delta_{O}^{B}$ and the subjective part $\Delta_{S}^{B}$, which is weighted by the average type. Finally, all of that only materializes with the probability of a buy recommendation, $\pi_{B}$.
@@
As the absolute value of $\Delta_{S}^{B}$ increases, the analysis becomes more complex because some types never accept recommendations. We consider the case $\Delta_{S}^{B}>2\Delta_{O}^{B}$ and reproduce the expression for $V(\mathcal{R})$ as follows: [^18] The argument for $\Delta_{S}^{B}<-2\Delta_{O}^{B}$ is analogous.
@@
$\underbrace{\pi^{B}F(\tilde{i})\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]\right]}_{\text{types accept all recommendations}}+\underbrace{(1-\pi^{B})(1-F(\tilde{i}))\left[\Delta_{O}^{D}-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]\right]}_{\text{types never accept recommendations}}.$ (8)
@@
In this case, the buy recommendation shifts probability from $(1,0)$ to $(0,1)$, which is unfavorable for receivers $i>0$, reducing their gains from accepting and possibly making them negative. Thus, for $\Delta_{S}^{B}>2\Delta_{O}^{B}$, only types below $\tilde{i}$ (a fraction $F(\tilde{i})$ of the receivers) accept all recommendations, which comes with probability $\pi^{B}$. In addition to the objective value of the recommendation ($\Delta_{O}^{B}$), these receivers obtain a subjective value. In (8), the subjective value is reflected in $-\Delta_{S}^{B}\mathbb{E}[i\mid i\leq\tilde{i}]$, the average effect of the shift in the relative probabilities of $(1,0)$ and $(0,1)$ on the value of accepting the recommendation for these buyers. The sign of this term depends on the properties of $F$ captured in $\mathbb{E}[i\mid i\leq\tilde{i}]$. It is positive if the negative contribution of types $i<0$ dominates the positive contribution of types $i>0$. The second component in (8), corresponding to the types who never accept recommendations, is arguably more surprising. Given their lack of alignment with the rest of the population, these types treat the recommendation not to buy as good news about the product and thus do not accept the recommendation. [^19] In detail, a don’t-buy recommendation (which happens with probability $1-\pi^{B}$) carries an adverse objective content (summarized by $\Delta_{O}^{D}$ in the term in squared brackets). However, for receivers with $i>0$ it also carries the good news that the product is more likely to be of the preferred version $(1,0)$ rather than $(0,1)$. Those with $i>\tilde{i}(>0)$ then buy the product despite the adverse objective content, valuing the beneficial subjective effect with $-\Delta_{S}^{D}\mathbb{E}[i\mid i\geq\tilde{i}]>0$ on average. Such receivers exist only if $\Delta_{S}^{B}$ is sufficiently large (and hence $\Delta_{S}^{D}$ is sufficiently negative relative to $\Delta_{O}^{D}$). Nonetheless, its informational content creates value for them.
@@
Going forward, we shall often invoke the following symmetry assumption on the preference distribution.
@@
We now define the quantity
$\beta:=F(1/2-R).$
Due to Assumption 1, it has a simple interpretation: Given a threshold $R$, the parameter $\beta$ captures the fraction of the population that is willing to recommend product versions $(1,0)$ or, equivalently (given the symmetry of the distribution), $(0,1)$. For a fixed type distribution $F$, the variables $\beta$ and $R$ are inversely related. We then obtain the following result:
@@
Suppose Assumption 1 holds. Then, all types accept all recommendations and
@@
To understand why all types accept all recommendations, the belief-updating decomposition in Section 4.1 is helpful. Its third component captures the potential shift of probabilities between controversial products, which induces a purely type-dependent payoff effect. Given the symmetry assumption, both controversial product versions lead to a buy recommendation with probability $\beta$. Hence, a recommendation does not shift probabilities between the controversial products. As the other two effects of a buy recommendation on payoffs are positive for all types, the buy recommendation increases expected payoffs for any type, yielding the result. [^20] An analogous argument applies for don’t-buy recommendations. Proposition 3 thus implies $V(R)=\pi^{B}\left[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]\right]$. Further, because of the symmetry of the population, $\mathbb{E}[i]=0$, so $V(R)=\pi^{B}\Delta_{O}^{B}$ as required.
@@ 5.2.1 Symmetric Type Distribution
By Lemma 2, an increase in $R$ reduces the probability of a buy recommendation, while it increases $\Delta_{O}^{B}$, the expected average payoff conditional on a buy recommendation. These effects exactly offset each other if and only if $\sigma=1$. Otherwise, one effect dominates the other. Based purely on uncontroversial products, a buy recommendation would be more likely than a don’t-buy recommendation when the good product is relatively more likely than the bad one ($\sigma>1$). It is thus worthwhile to increase the expected average payoff conditional on a buy recommendation by increasing $R$ at the expense of the probability of receiving such a recommendation. Put differently, the designer chooses between revealing whether the product is good or whether it is bad. He moves the threshold in the direction which yields full revelation with higher probability, which hinges only on the odds ratio of the uncontroversial product versions.
@@ 5.2.2 Asymmetric Type Distribution
We can now analyze the extent to which the result in Proposition 5 that, generically, extreme thresholds are desirable is an artifact of the symmetry assumption. This is challenging because, contrary to the symmetric case, the set of types that accept all recommendations depends on the threshold $R$. Reflecting this complication, there is no direct counterpart to Proposition 5. Nevertheless, the following two results provide some insights into the optimal recommendation system in the case of an asymmetrically distributed population.
@@
As illustrated in Figure 2, Proposition 6 provides sufficient conditions for interior solutions. It shows that an interior optimum exists when (i) controversial products are likely ($Q$ is large) and the population is very asymmetric ($a$ is far away from $1$) or otherwise when (ii) the uncontroversial product versions are similarly likely ($\sigma$ is close to $1$). [^25] The figure might suggest that the parameter region where $R^{\ast}\rightarrow 1$ is in some sense larger than the area where $R^{\ast}\rightarrow 0$. This is an artefact of the otherwise quite convenient parameterization of the preference distribution but does not have any significant economic meaning. The result starkly contrasts the symmetric case, where an interior solution is only possible in the knife-edge case $\sigma=1$. In that case, recommendations do not reveal any information about the relative likelihood of the controversial products beyond the prior, so the subjective payoff effect cancels out and the optimal threshold depends only on the odds ratio $\sigma$ of the uncontroversial products. In contrast, the average subjective payoff effect does not cancel out in the asymmetric case. Thus, the choice of the optimal threshold no longer only depends on $\sigma$ but also on the asymmetry parameter $a$ and the prevalence of controversial products $Q$. We explore under which circumstances the conditions given in Proposition 6 are necessary as well as sufficient:
@@
The proof essentially proceeds in two steps. First, all types accept all recommendations when $Q$ is sufficiently small, since then the subjective content in the recommendation vanishes.262626 $Q\rightarrow 0$ implies that $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$, ensuring that all types accept all recommendations by Corollary 1. The key step is that $\Delta_{S}^{B}\rightarrow 0$ as $Q\rightarrow 0$, while $\Delta_{O}^{B}\rightarrow 1-q_{H}>0$ when $Q\rightarrow 0$. The proof shows that, even for larger values of $Q$, $|\Delta_{S}^{B}|\leq 2\Delta_{O}^{B}$ often holds, explaining why the conclusion of Proposition 7 holds not only in the limit. This simplifies $V(R)$ by eliminating the case distinctions. Leveraging this, we can prove the quasiconcavity of the function, ensuring that the sufficient conditions in Proposition 6 are necessary, too.
@@
The requirement that $Q\rightarrow 0$ is much stronger than necessary for this result. Even for higher values of $Q$, Figure 2 quite accurately separates the parameter range under which an interior solution exists from the area where it does not. [^27] The figure is drawn for $Q=0.1$, in which case the if-and-only-if statement is correct for $\sigma\in[1/60,60]$ and $a\in[1/10000,10000]$. In a nutshell, the result shows that an interior solution exists if (i) controversial products are sufficiently prevalent, (ii) the population asymmetry is substantial, and (iii) the odds ratio of the uncontroversial products is sufficiently close to 1. To understand this intuitively and to contrast it with the symmetric case, suppose that $\sigma>1$. Hence, the value of the recommendation system is increasing in $R$ in the symmetric case, as the gain in the expected payoff conditional on receiving a buy recommendation outweighs the loss from the lower probability of a buy recommendation. This follows because the subjective effect in $V(R)=\pi^{B}[\Delta_{O}^{B}-\Delta_{S}^{B}\mathbb{E}[i]]$ cancels out with symmetry. With an asymmetric population, the term $-\Delta_{S}^{B}\mathbb{E}[i]$ does not vanish. To illustrate this, consider a very asymmetric preference distribution such as $F(x)=(x+1/2)^{10}$, for which the probability of a type $i\leq 0$ is less than 0.1%. Then, $\mathbb{E}[i]>0$ and, by Corollary 2, $\Delta_{S}^{B}<0$ so that $-\Delta_{S}^{B}\mathbb{E}[i]>0$. Hence, the subjective effect is positive on average. Intuitively, though this population is distributed very asymmetrically around zero, it is essentially homogeneous in the sense that almost all types agree on the ranking of the controversial products: If such a product results in a buy recommendation, it almost surely has quality vector (1,0), as almost all types prefer this to $(0,1)$. As such, these products are not that controversial anymore, and even the subjective effect is “fairly objective.” Now, as $R$ increases, the probability of a buy recommendation still falls, and the “purely objective” payoff part $\Delta_{O}^{B}$ increases. However, for $R$ high enough, this “fairly objective” payoff part $-\Delta_{S}^{B}\mathbb{E}[i]$ also decreases, as an extreme threshold $R$ impedes the ability to distinguish between a good product and a widely preferred controversial product. While consumers are homogeneous regarding which controversial product version they prefer, they are still heterogeneous with respect to the intensity with which they do so. Hence, disentangling the good and the preferred controversial product is worthwhile. Therefore, an interior optimal threshold results.
@@ 5.3.1 Preference Heterogeneity
In this section, we analyze how changes in the (exogenously given) preference distribution affect the value of the recommendation system. [^28] When we consider distinct sender and receiver distributions in Section 6.1, we will build further on this and take a design perspective instead of the comparative statics approach here. Under Assumption 1, we can analyze increases in the polarization of the distribution of preferences by keeping the threshold $R$ fixed and varying $\beta$ [^29] For $R=1/2$, we must have $\beta=1/2$ because the symmetry of $F$ requires $F(0)=1/2$, so this exercise is only meaningful for thresholds $R\neq 1/2$. Further, for $R>1/2$ we must have $\beta\leq 1/2$ because $\beta=F(1/2-R)\leq F(0)=1/2$. Accordingly, for $R<1/2$ we must have $\beta\geq 1/2$. This corresponds to a mean-preserving spread of $F$ (see Figure 3), coinciding with an increase in $\beta$ for $R>1/2$ and a decrease for $R<1/2$. We obtain the following result.